Please Help Me!!!!
I need to find the equation of the line passing through the point
(2,7) with a slope of m=2
A)y=1/2x-3/2
B)y=2x-12
C)y=1/2x+6
D)y=2x+3
Please show your work, so that I can learn how this is done...thanks

Answer by elima(1433)

(2,7) with a slope of m=2
the formula would be;
(y-Y1)=m(x-X1)
m=slope
(X1,Y1)= the point given;
(y-7)=2(x-2)
y-7=2x-4
y=2x-4+7
y=2x+3
So the answer is D.
=)